# code-for-first-row = , code-for-last-row = , code-for-first-col = , code-for-last-col =

### v1

@@ Abstract
Posets are discrete mathematical structures which are ubiquitous in a broad range of data analysis and machine learning applications. Research connecting posets to the data science domain has been ongoing for many years. In this paper, a comprehensive review of a wide range of studies on data analysis amd machine learning using posets are examined in terms of their theory, algorithms and applications. In addition, the applied lattice theory domain of formal concept analysis will also be highlighted in terms of its machine learning applications.
@@ 1 Introduction
Data analysis plays a pivotal role in today’s data-driven world. Machine learning is used to automate the data analysis process, and enables the the extraction of valuable insights by making accurate predictions from large complex datasets. Data often have complex structure which can usually be endowed with a natural order. Such complex data can be derived from text, chemical structures, biological structures, images, videos and so on. Mathematical techniques and concepts serve as powerful tools in the development of new algorithms and methods for machine learning and data analysis. For some time now, many studies based on the use of partially ordered sets(posets) for machine learning and data analysis have emerged. Furthermore, over this time the literature on these topics has significantly expanded in volume, scope, and the range of its applications in various fields. However, no recent surveys exist to collect and organize this knowledge thereby impeding the ability of researchers and machine learning scientists alike to utilize it for various data science or artificial intelligence tasks. In order to address this gap, we present an up-to-date overview with an emphasis on the most prominent and currently relevant works on data analysis and machine learning methods in which the role of posets has been established. By covering earlier works as well as more recent advances, our survey aims to provide researchers, applied mathematicians, data scientists, statisticians, social scientists and practitioners new to the field as well as all those aiming at keeping pace with innovation in this interesting, growing and promising research field with a solid understanding of the different conceptual and methodological approaches which incorporates posets into data analytics or machine learning applications.
The rest of this survey is organized as follows. Section 2 introduces and outlines the key concepts on partially ordered sets and the related field of lattice theory. Section 3 presents a summary of various research studies on machine learning and deep learning with respect to poset theory, lattice theory and formal concept analysis. Section 4 focuses on cluster analysis using posets and formal concept analysis methods. Sections 5-7 presents a summary on data analysis from a poset theoretical perspective covering its multidimensional, exploratory and descriptive aspects respectively. Section 8 presents a collection of applications across various domains including software packages, datasets and algorithms for this purpose. Section 9 discusses the current challenges and suggests future directions. Section 10 summarizes the paper.
@@ 2 Basics of Partial Order Theory
    antisymmetric: $x,y\in X$ if $x\preccurlyeq y$ and $y\preccurlyeq x\implies x=y,$
@@
    transitive: $x,y,z\in X$ if $x\preccurlyeq y$ and $y\preccurlyeq z\implies x\preccurlyeq z.$
@@
  • Maximal elements or objects of the poset $P$ are the set of elements or objects of $X$ for which no other object or element $y$ satisfies the relation $y>x.$ If $x$ is the only maximal object or element then it is referred to as the ”greatest” object or element. In a totally ordered set, the terms maximal element or object and greatest element or object coincide.
  • Minimal elements or objects of the poset $P$ are the set of elements or objects of $X$, for which no other object or element $y$ exist such that $y<x$. If $x$ is the only minimal object or element then $x$ is referred to as the “least” object or element. In a totally ordered set, the terms minimal element or object and least element or object coincide.
  • Chain is a subset C of X, where any object or eleement is mutually comparable with all other objects or elements of C. That is, a chain is a poset such that no element or object can be added between any two of its elements without losing the property of being totally ordered.
  • Antichain is a subset $C^{\prime}$ of X, where each object or element of $C^{\prime}$ is mutually incomparable with all other objects or elements in $C^{\prime}$. That is, all the objects or elements of the set $C^{\prime}$ are pairwise incomparable elements and as such they can never have the property of being totally ordered.
  • A cover relation for ${\mathcal{P}}=(X,\preccurlyeq)$ is the set of pairs $(x,y)$ such that $y\in X$ covers $x$ whenever there exist no element $z\in X$ such that $x<z<y.$
Partially ordered sets can often be graphically visualized using Hasse diagrams[62]. A Hasse diagram can be derived from a directed acyclic graph, where the vertices are representing the objects and a line relates object $x$ with $y$ whenever $x<y$. In the case of a transitivity relation whenever $x<y$ and $y<z$ then it suffices to draw a line only for $x<y$ and $y<z$ and not for $x<z$. Hasse diagrams can be viewed as a powerful tool for unifying ideas and concepts. Hasse diagrams reveal an intricate network of comparabilities and incomparabilities, maximal and minimal elements. In addition, Hasse diagrams are also characterized with a structure that comprises levels, chains and antichains.
@@
Finite posets can also be represented by square matrices which comprise: incidence [262, 262], cover [263], and mutual ranking probability matrices[264]. The mutual ranking probability matrices are a class of matrices which comprise and convey information on the dominance among statistical units All the three main types of poset matrices are useful in conveying essential information on the structure of the order relation from different perspectives.
@@
Posets can also be subdivided into variour classes such as series-parallel posets, semantic posets and ranking posets etc. Series-parallel posets are characterized by the N-free partial orders. Series-parallel partial orders have been applied machine learning of event sequencing in time series data [134]. Semantic posets which are used for compositional generalization in semanti parsing [105]. The ranking poset $\mathfrak{M}_{n}$[129] is the poset in which the elements are all possible cosets $\mathfrak{G}_{\lambda}\pi$, where $\lambda$ is an ordered partition of $n$ and $\pi\in\mathfrak{G}_{n}.$ The ranking poset may be useful in model selection and validation in machine learning.
@@
Posets are natural models for many statistical applications [219, 165]. Partial orders are the natural mathematical structure for comparing multivariate data that lack a natural order [143]. For instance, multivariate data on colours has no natural ordering and so the rank features in colour spaces are based on partial orders and this approach is generalizable to similar contexts [276]. It also plays a key role in decision making in environmental science and chemistry [33, 107] etc.
@@
The study of partially ordered sets and lattices has been covered by an extensive amount of mathematical literature. Birkhoffs (1967) [258] and Grdtzer’s (1978)[259] books on lattice theory are considered to be classics. Lattices[181] are a broad structured class of posets. Lattice theory has shown its ability to contribute to solving problems in a wide variety of applications. Formally, a lattice [104] is a poset, in which every pair of elements has both a least upper bound and a greatest lower bound. In other words, it is a structure with two binary operations: join and meet. The following key sub-definitions summarize a lattice structure. A meet semilattice is a poset for which any two elements a and b have a greatest lower bound denoted $a\wedge b$. The greatest lower bound of a and b is the largest element that is still less than both of them. In a lattice, the greatest lower bound must be unique. The greatest lower bound of a and b is also called the meet or infimum of $a$ and $b$. A join semilattice is a poset for which any two elements $a$ and $b$ have a least upper bound, denoted $a\vee b$. The least upper bound of $a$ and $b$ is the smallest element that is still greater than both. In a lattice, the least upper bound must be unique. The least upper bound of $a$ and $b$ is also called the join or supremum of $a$ and $b$. If a poset is both a meet semilattice and a join semilattice, then the poset is also a lattice.
@@ 3.1 Overview
Machine learning generally involves the creation of knowledge from a training dataset by using models and algorithms to learn from data in order to make predictions, find pattrens, classify data or automate decision making processes. Machine learning has progressed significantly over the past two decades, and this has been driven by the development of new learning algorithms, theory, and high performance computing architecture [290]. There are many different kinds of machine learning algorithms. The most well-known ones are supervised, unsupervised, semi-supervised, and reinforcement learning. The goal of machine learning algorithm is to optimize the performance of a system when handling new instances of data through user defined programming logic for a given environment [301]. In order to build more efficient machine learning algorithms or models so as to improve on their performance and accuracy in most predictive task, various ensemble learning methods have been developed. Ensemble learning[291] is an approach which aggregates the predictions of two or more models fitted to the same data in order to minimize error. In the area of supervised learning, the most popular ensemble machine learning techniques are bagging, boosting and stacking, which use multiple learning algorithms or models to produce one optimal predictive model. Ensemble methods can broadly be categorized into sequential ensemble techniques and parallel ensemble techniques. Bagging aims to reduce variance by adopting parallel ensemble learning, boosting aims to reduce bias by adopting sequencial ensemble learning , and stacking aims to improve prediction accuracy by adopting parallel ensemble learning. Sequential ensemble techniques generate base learners in a sequence that are characterized by the dependence between the base learners. On the other hand the parallel ensemble technique base learners are generated in parallel in such a way that independence is enforced between the base learners. Dagging(Disjoint samples aggregating) is another useful type of parallel ensemble technique that was introduced in [300]. Dagging is similar to bagging, but instead of bootstrapping, it uses stratified sampling by creating a number of disjoint groups and stratified data from the original learning data set, with each considered as a subset of learning. In unsupervised learning, ensemble clustering aims at combining the outputs of several clustering algorithms to form a single clustering structure (crisp or fuzzy partition, hierarchy) [302, 303].
@@
In recent years, deep learning[292] has been the most popular computational approach in the field of machine learning and has unique advantages when dealing with high-dimensional nonlinear problems. The performance of the machine learning algorithm relies heavily on the representation and finding a good representation can facilitate discovering structure in the input data by the learning algorithm. Deep learning as a model is a specific type of representation learning[283] that uses layered algorithms known as artificial neural networks, which attempts to mimic the human brain through a combination of data inputs, weights, and bias, to learn representations of data. It has succesfully been used in a wide range of disciplines such as disciplines, including cybersecurity, natural language processing, visual recognition, machine translation, robotics, recommendation systems etc. There are several types of deep learning algorithms. These are : convolutional Neural Network, Long Short Term Memory, GraphNeural Networks etc. Traditionally, machine-learning and deep learning algorithms have been used on data represented in euclidean space such as image data which can be represented as a regular grid of pixel values. On the other hand, graph data cannot be represented on euclidean space [307]. Graph Neural Networks (GNNs) are a class of deep learning methods designed to perform inference on data described by graphs which are non-euclidean (non-metric) structured. Posets are special classes of directed graphs. There is also a growing use of graph neural networks for the modeling and analysis of partially ordered data [99, 83, 105, 304, 306]. More recently, the problem of machine learning on meet/join lattices and posets was extensively studied in [180]. The main focus of the study was on developing methods that are based on generalized convolutions and sparse Fourier transforms algorithms (SFTs) that are capable of learning set functions. The first lattice-based machine learning models were introduced by V.K. Finn[295, 293, 294, 296]. It was based on a closure systems JSM-method1 of automated hypothesis generation. In this model positive hypotheses are searched among intersections(similarity as meet operations) of positive example descriptions (object intents), likewise for negative hypotheses [297]. The JSM method also identifies data patterns by means of induction and it can also be used as a logical rule-based classification method formulated in terms of formal concepts [298]. Machine learning has been studied from many perspectives in the context of poset theory and lattice theory. These include : Machine Learning Performnce Comparison[25, 129], Multilabel Classification [76], Ensemble Classification[129], Sequential Clasification [134, 166, 166, 165, 172]; Natural Language Processing [83, 132, 105], Deep Unsupervised Learning [20], Semi-Supervised Learning[99], Learning to Rank[270], Time Series Modelling [288, 134] .
3.2 Comparing the performance of machine learning algorithms using posets
@@
Performance comparison of different learning algorithms is a very common task in the field of machine learning. The evaluation of several machine learning algorithms based on the concept of depth functions was recently introduced in [25]. Depth functions [218] were originally formulated to extend the univariate notions of median, quantiles, ranks, signs, and order statistics to the setting of multivariate data. Blocher & Schollmeyer [25] uses an adaptation of the well-known simplicial depth and then applied it to the set of all partial orders referred to as the union-free generic (ufg) depth. The purpose is to develop a framework to show how samples of poset-valued random variables can be descriptively analyzed by utilizing a generalized concept of data depth based on the concept of depth functions. A key property of the ufg depth is that it considers partial orders as a whole, and not limited to pairwise comparisons since as the procedure for computing the depth of a partial order cannot be broken down using simple sum statistics, but requires the partial order as a holistic entity. The ufg depth for a comparison of machine learning algorithms is based on multidimensional performance measures. It works by comparing the various algorithms not with respect to one unidimensional criterion such as balanced accuracy, but instead it considers a whole set of performance measures.

### v2

@@ Abstract
Posets are discrete mathematical structures which are ubiquitous in a broad range of data analysis and machine learning applications. Research connecting posets to the data science domain has been ongoing for many years. In this paper, a comprehensive review of a wide range of studies on data analysis and machine learning using posets are examined in terms of their theory, algorithms and applications. In addition, the applied lattice theory domain of formal concept analysis will also be highlighted in terms of its machine learning applications.
@@ 1 Introduction
Data analysis plays a pivotal role in today’s data-driven world. Machine learning is used to automate the data analysis process, and enables the extraction of valuable insights by making accurate predictions from large complex datasets. Data often have complex structure which can usually be endowed with a natural order. Such complex data can be derived from text, social and behavioral sciences, chemical structures, biological structures, images, videos, governmental and business settings, and so forth. Mathematical techniques and concepts serve as powerful tools in the development of new algorithms and methods for machine learning and data analysis. For some time now, many studies based on the use of partially ordered sets(posets) for machine learning and data analysis have emerged. Furthermore, over this time the literature on these topics has significantly expanded in volume, scope, and the range of its applications in various fields. However, no recent surveys exist to collect and organize this knowledge, thereby impeding the ability of researchers and machine learning scientists alike to utilize it for various data science or artificial intelligence tasks. In order to address this gap, we present an up-to-date overview with an emphasis on the most prominent and currently relevant works on data analysis and machine learning methods in which the role of posets has been established. By covering earlier works as well as more recent advances, this survey aims to provide researchers, applied mathematicians, data scientists, statisticians, social scientists and practitioners new to the field as well as all those aiming at keeping pace with innovations in this interesting, growing and promising research field with a solid understanding of the different conceptual and methodological approaches which incorporates posets into data analytics or machine learning applications.
The rest of this survey is organized as follows. Section 2 introduces and outlines the key concepts on poset theory and lattice theory. Section 3 presents a summary of research studies on machine learning and deep learning with respect to poset theory, lattice theory and formal concept analysis. Section 4 focuses on cluster analysis using posets and formal concept analysis methods. Sections 5-7 presents a summary on data analysis from a poset theory perspective covering its multidimensional, exploratory and descriptive aspects respectively. Section 8 presents a collection of applications across various domains including software packages, datasets and algorithms selected for this purpose. Section 9 discusses possible future research directions. Section 10 summarizes the paper.
@@ 2 Basics of Partial Order Theory
    antisymmetric i.e. $x,y\in X$ if $x\preccurlyeq y$ and $y\preccurlyeq x\implies x=y;$
@@
    transitive i.e. $x,y,z\in X$ if $x\preccurlyeq y$ and $y\preccurlyeq z\implies x\preccurlyeq z.$
@@
  • Maximal elements (or objects) of the poset $P$ are the set of elements $x\in X$ in which no other element $y\in X$ satisfies the relation $y>x.$ If $x$ is the only maximal element then it is referred to as the “greatest” element. In a totally ordered set, the terms maximal element and greatest element coincide.
  • Minimal elements (or objects) of the poset $P$ are the set of elements $x\in X$ in which no other element $y\in X$ exist such that $y<x$. If $x$ is the only minimal element, then $x$ is referred to as the “least” element. In a totally ordered set, the terms minimal element and least element coincide.
  • Chain is a subset C of X, in which any object (or element) is mutually comparable with all other elements of C. That is, a chain C is a poset such that no element can be added between any two of its elements without losing the property of being totally ordered.
  • Antichain is a subset $C^{\prime}$ of X, in which each object (or element) of $C^{\prime}$ is mutually incomparable with all other elements in $C^{\prime}$. That is, all the elements of the set $C^{\prime}$ are pairwise incomparable and as such they can never have the property of being totally ordered.
  • A cover relation for ${\mathcal{P}}=(X,\preccurlyeq)$ is the set of pairs $(x,y)$ such that $x,y\in X,$ and $y$ covers $x$ whenever there exist no element $z\in X$ such that $x\prec z\prec y.$
Partially ordered sets are graphically visualized using Hasse diagrams. A Hasse diagram can be derived from a directed acyclic graph, where the vertices are representing the objects and a line relates object $x$ with $y$ whenever $x\prec y$. In the case of a transitivity relation whenever $x\prec y$ and $y\prec z$ then it suffices to draw a line only for $x\prec y$ and $y\prec z$ and not for $x\prec z$. Hasse diagrams can be viewed as a powerful tool for unifying ideas and concepts. Hasse diagrams reveal an intricate network of comparabilities and incomparabilities, maximal and minimal elements. In addition, Hasse diagrams are also characterized with a structure that comprises levels, chains and antichains.
@@
Finite posets can also be represented by square matrices which comprise: incidence [98, 223], cover [7], and mutual ranking probability matrices[131]. The mutual ranking probability matrices are a class of matrices which comprise and convey information on the dominance among statistical units[131]. All the three main types of poset matrices are useful in conveying essential information on the structure of the order relation from different perspectives. Posets can also be subdivided into variour classes such as series-parallel posets, semantic posets, and ranking posets etc. Series-parallel posets are characterized by the N-free partial order structure. Series-parallel partial orders have been applied to machine learning of event sequencing in time series data [212]. Semantic posets are used for compositional generalization in semantic parsing [150]. The ranking poset $\mathfrak{M}_{n}$[204] is the poset in which the elements are all possible cosets $\mathfrak{G}_{\lambda}\pi$, where $\lambda$ is an ordered partition of $n$ and $\pi\in\mathfrak{G}_{n}.$ The ranking poset is useful in model selection and validation in machine learning.
@@
Posets are natural models for many statistical applications [38, 272, 235, 270]. Partial orders are also the natural mathematical structure for comparing multivariate data that lacks a natural order [234]. For instance, multivariate data on colours has no natural ordering and therefore the rank features in colour spaces can be modelled as partial orderings, and this approach is generalizable in similar contexts [257]. It also plays a key role in decision making in environmental informatics and computational chemistry [52, 156, 192, 194, 193, 55, 63], social sciences[28, 67] etc. Moreover, in data analysis a typical binary relation is defined for a set of objects and a set of attributes. For instance, objects are transactions in a supermarket, attributes are items in the supermarket, and the relation consists of pairs (transaction, item) such that the item occurs in the transaction[203].
@@
The study of partially ordered sets and lattices has been covered by a significant number of mathematical publications. Birkhoffs (1967) [37] $\ddot{\text{a}}$tzer’s (1978)[149] books on lattice theory are considered to be classics. Lattices[297] are a broad structured class of posets. Lattice theory has shown its ability to contribute to solving problems in a wide variety of applications. Formally, a lattice[147] is a poset, in which every pair of elements has both a least upper bound and a greatest lower bound. In other words, it is a structure with two binary operations: join and meet. The following sub-definitions summarize a lattice structure. A meet semilattice is a poset for which any two elements a and b have a greatest lower bound denoted $a\wedge b$. The greatest lower bound of a and b is the largest element that is still less than both of them. In a lattice, the greatest lower bound must be unique. The greatest lower bound of a and b is also called the meet or infimum of $a$ and $b$. A join semilattice is a poset for which any two elements $a$ and $b$ have a least upper bound, denoted $a\vee b$. The least upper bound of $a$ and $b$ is the smallest element that is still greater than both. In a lattice, the least upper bound must be unique. The least upper bound of $a$ and $b$ is also called the join or supremum of $a$ and $b$. If a poset is both a meet semilattice and a join semilattice, then the poset is also a lattice.
@@ 3.1 Overview
In general, machine learning involves the creation of knowledge from a training dataset by using models and algorithms to learn from data in order to make predictions, find patterns, classify data and automate decision making processes. Machine learning has progressed significantly over the past two decades, and this has been driven by the development of new learning algorithms, theory, and high performance computing architecture [145]. There are many different kinds of machine learning algorithms. The most well-known ones are supervised, unsupervised, semi-supervised, and reinforcement learning. Supervised learning is useful in scenarios for which some relationship between input and output labelled data has been established, unsupervised learning is effective for uncovering hidden patterns in unlabelled data, and semi-supervised learning utilizes a combination of labelled and unlabelled data to train models. The goal of any machine learning algorithm is to optimize the performance of a system when handling new instances of data through user defined programming logic in a given environment [29]. In order to build more efficient machine learning algorithms or models so as to improve on their performance and accuracy in most predictive task, various ensemble learning methods have been developed. Ensemble learning[246] is an approach which aggregates the predictions of two or more models fitted to the same data in order to minimize error. In the area of supervised learning, the most popular ensemble machine learning techniques are bagging, boosting, and stacking, which use multiple learning algorithms or models to produce one optimal predictive model. Ensemble methods can broadly be categorized into sequential ensemble techniques and parallel ensemble techniques. Bagging(bootstrap aggregating) aims to reduce variance by adopting parallel ensemble learning on homogenous weak learners, boosting aims to reduce bias by adopting sequencial ensemble learning on homogenous weak learners, and stacking aims to improve prediction accuracy by adopting parallel ensemble learning on heterogeneous weak learners. Sequential ensemble techniques generate base learners in a sequence that are characterized by the dependence between the base learners. On the other hand the parallel ensemble technique base learners are generated in parallel, and in such a way that independence is enforced between the base learners. Dagging(Disjoint samples aggregating) is another useful type of parallel ensemble technique that was introduced in [280]. Dagging is similar to bagging, but instead of bootstrapping, it uses stratified sampling by creating a number of disjoint groups and stratified data from the original learning data set, with each considered as a subset of learning. In unsupervised learning, ensemble clustering[306, 19] aims at combining the outputs of several clustering algorithms to form a single clustering structure (crisp or fuzzy partition, hierarchy).
@@
In recent years, deep learning[114] has been the most popular computational approach in the field of machine learning and has unique advantages when dealing with high-dimensional nonlinear problems. The performance of machine learning algorithms relies heavily on the representation, and finding a good representation can facilitate the discovery of structure in the input data by the learning algorithm. Deep learning as a model is a specific type of representation learning[33] that uses layered algorithms known as artificial neural networks, which attempts to mimic the human brain through a combination of data inputs, weights, and biases, to learn representations of data. It has succesfully been used in a wide range of disciplines such as cybersecurity, natural language processing, visual recognition, machine translation, robotics, recommendation systems etc. There are several types of deep learning algorithms. These are: Convolutional Neural Network(CNN), Long Short Term Memory, Graph Neural Networks etc. Traditionally, machine-learning and deep learning algorithms have been used on data represented in Euclidean space such as image data which can be represented as a regular grid of pixel values. On the other hand, graph data cannot be represented on Euclidean space [20]. Graph Neural Networks (GNNs)[299] are a class of deep learning methods designed to perform inference on data described by graphs which are non-Euclidean (non-metric) structured. They are highly influenced by CNN which learns features by inspecting neighboring pixels(nodes in GNN) in three dimensional image data for classification and object recognition purposes. Posets are special classes of directed graphs. There is also a growing use of graph neural networks for the modeling and analysis of partially ordered data [132, 116, 150, 295, 304]. More recently, the problem of machine learning on meet/join lattices and posets was extensively studied in [296]. The main focus of the study was on developing methods that are based on generalized convolutions and sparse Fourier transforms algorithms which are capable of learning set functions. The first lattice-based machine learning models were introduced by V.K. Finn[137, 136, 135]. It was based on a closure system which uses the JSM(John Stuart Mille)-method of automated hypothesis generation. In this model, positive hypotheses are searched among intersections(similarity as meet operations) of positive example descriptions (object intents), likewise for negative hypotheses [21]. The JSM method also identifies data patterns by means of induction and it can also be used as a logical rule-based classification method formulated in terms of formal concepts [200]. Machine learning has been studied from many perspectives in the context of poset theory and lattice theory. These include : machine learning performnce comparison[39, 204], multilabel classification [109], ensemble classification[204], sequential clasification [212, 273, 272, 278]; natural language processing [116, 210, 150], deep unsupervised learning [31], semi-supervised learning[132], learning to rank[96], time series modelling [228, 212], ensemble clustering [109], model selection[270], topological deep learning [155], attention-based neural networks [154].
3.2 Performance Comparison and Model Selection
@@
Performance comparison of different learning algorithms is a very common task in the field of machine learning. The evaluation of several machine learning algorithms based on the concept of depth functions was recently introduced in [39]. Depth functions[312] were originally formulated to extend the univariate notions of median, quantiles, ranks, signs, and order statistics to the setting of multivariate data. Blocher & Schollmeyer [39] uses an adaptation of the well-known simplicial depth, and then applied it to the set of all partial orders referred to as the union-free generic (ufg) depth. The purpose is to develop a framework to show how samples of poset-valued random variables can be descriptively analyzed by utilizing a generalized concept of data depth based on the concept of depth functions. A key property of the ufg depth is that it considers partial orders as a whole, and not limited to pairwise comparisons, since as the procedure for computing the depth of a partial order cannot be broken down using simple sum statistics, but requires the partial order as a holistic entity. The ufg depth for a comparison of machine learning algorithms is based on multidimensional performance measures. It works by comparing the various algorithms not with respect to one unidimensional criterion such as balanced accuracy, but instead it considers a whole set of performance measures. In Taeb et al.[270], a methodology is presented for organizing classes of models as partially ordered sets in order to address the challenges of model structures which are not characterized by an underlying Boolean logical structure, and which is a prerequisite to formalize and control the false-positive error.